Find the two x- intercept of \\ (f(x)=x^{2}-3x+2\\)


1
Expert's answer
2019-09-09T11:42:37-0400

Solution. For x- intercept of f(x) coordinate y=0 (f(x)=0). Find the solution to the equation


"x^2-3x+2=0""D=(-3)^2-4\times1\times2=9-8=1"

Hence, we have two solutions


"x_1=\frac {3-\sqrt 1} {2}=1, x_2=\frac {3+\sqrt 1} {2}=2."

Therefore get two points


"(1,0), (2,0)."

Answer. (1,0), (2,0).